OpenStudy (anonymous):

how do you turn this (vertex form) to standard form? y= -(x-1)^2+5

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

Follow through with (x-1)^2 and write out the full expression

OpenStudy (anonymous):

would it be -(x-1)(x-1) or how does the negative work?

OpenStudy (anonymous):

@amistre64 please help

OpenStudy (amistre64):

the negative will effect the product, changes signs is all

OpenStudy (amistre64):

-(4)^2 = -16

OpenStudy (amistre64):

-(x-1)^2 -(x^2-2x+1) -x^2 +2x -1

OpenStudy (amistre64):

then add the 5 in :)

OpenStudy (anonymous):

so -x^2-2x+1 to start with? then +5 to the +1 ?

OpenStudy (anonymous):

oops -1*

OpenStudy (amistre64):

-(x^2-2x+1) , swap out ALL the signs

OpenStudy (amistre64):

your distributing a -1 thru the paranthsis

OpenStudy (anonymous):

do you distribute a negative 1 even to the second set of parenthesis?

OpenStudy (amistre64):

youve already taken care of the second set ... by "foiling" or however you want to describe it

OpenStudy (amistre64):

\[-(x-1)^2\] \[-(x-1)(x-1)\] \[-(x^2-2x+1)\] \[-x^2--2x-+1\] \[-x^2+2x-1\]

OpenStudy (anonymous):

basically I am trying to find the zero's of the equation

OpenStudy (amistre64):

finding the zeros does not need to be determined from the standard form do you have to convert this to standard form, or just find the zeroes?

OpenStudy (anonymous):

just find the zero's

OpenStudy (amistre64):

then let y=0 \[0= -(x-1)^2+5\] \[(x-1)^2=5\] \[x-1=\pm\sqrt{5}\] \[x=1\pm\sqrt{5}\]

OpenStudy (anonymous):

thanks you helped out a lot and answered all my questions! :)

OpenStudy (amistre64):

:) good luck